# Overcomplete Tensor Decomposition via Koszul–Young Flattenings

### v1

@@ Appendix B Commuting Extensions
    Let $q$ be the largest odd integer with $q\leq m$, and let $p=(q-1)/2$.
@@
The rank-1 matrices used in our construction will each be supported on a rectangle, either with all nonzero values equal to $1$ or all nonzero values equal to $-1$. One possible construction is to take the rectangle $[n]\times[n]$ and subtract off a collection of rectangles that partition $([n]\times[n])\setminus\{(1,1),(2,2),\ldots,(n,n)\}$. The number of rectangles in the partition can range from $2(n-1)$ to $n^{2}-n$, so this covers all the $r$ values ranging from $12(n-1)+$ 1 to $n^{2}-n+1$. An alternative construction is to, for some $k\in[n]$, take the rectangle $[k]\times[k]$ and subtract off a collection of $2(k-1)$ rectangles that partition $([k]\times[k])\setminus\{(1,1),(2,2),\ldots,(k,k)\}$, and then add $n-k$ additional $1\times 1$ rectangles on the diagonal. This uses $11+2(k-1)+(n-k)=n+k-$ 1 rectangles in total, so by varying $k$ this covers all the $r$ values ranging from $n$ to $12n-$ 1. ∎

### v2

@@ Appendix B Commuting Extensions
    Let $q$ be the largest odd integer with $q\leq m$, and let p=(q1)/2p=(q-1)/2.
@@
The rank-1 matrices used in our construction will each be supported on a rectangle, either with all nonzero values equal to $1$ or all nonzero values equal to 1-1. One possible construction is to take the rectangle $n]\times[n]$ and subtract off a collection of rectangles that partition $([n]\times[n])\setminus\{(1,1),(2,2),\ldots,(n,n)\}$. The number of rectangles in the partition can range from 2(n1)2(n-1) to $n^{2}-n$, so this covers all the $r$ values ranging from 2(n1)+12(n-1)+1 to $n^{2}-n+1$. An alternative construction is to, for some $k\in[n]$, take the rectangle $k]\times[k]$ and subtract off a collection of 2(k1)2(k-1) rectangles that partition $([k]\times[k])\setminus\{(1,1),(2,2),\ldots,(k,k)\}$, and then add nkn-k additional $1\times 1$ rectangles on the diagonal. This uses 1+2(k1)+(nk)=n+k11+2(k-1)+(n-k)=n+k-1 rectangles in total, so by varying $k$ this covers all the $r$ values ranging from $n$ to 2n12n-1. ∎